The half-life of a radioactive substance is the time it takes for half of the substance to decay. The half-life of carbon-14 is 5700 years.
a. Write an exponential function to model the decay of a 470-mg sample.
b. Explain what each value in the function model represents.
c. To the nearest hundredth, find the amount of carbon-14 remaining after 2619 years. Explain how you found this amount.

Answer by ankor@dixie-net.com(22740)

The half-life of a radioactive substance is the time it takes for half of the substance to decay.
The half-life of carbon-14 is 5700 years.
:
a. Write an exponential function to model the decay of a 470-mg sample.
A = 470*2^(-t/5700)
:
b. Explain what each value in the function model represents.
A = resulting amt after t yrs
470 = initial amt
5700 = half life of the substance
t = time of decay for the substance
:
c. To the nearest hundredth, find the amount of carbon-14 remaining after 2619 years.
Explain how you found this amount.
A = 470*2^(-2619/5700)
A = 470*2^-.45947
A = 470 * .727
A = 341.81 mg after 2619 yrs
:
You should be able to explain what we did here